# StefaLand: An Efficient Geoscience Foundation Model That Improves Dynamic Land-Surface Predictions

### v2

@@ 3 Experiments
We tested the value of foundation model pretraining on 5 datasets and 6 experiments, including, streamflow on the CAMELS dataset on USA, CAMELS streamflow prediction with hybrid model, global streamflow, global in-situ soil moisture, global soil properties, and landslide susceptibility in Oregon, USA. For all experiments, hyperparameters were tuned with Ray Tune and kept consistent across model configurations within each experimental case (e.g., CAMELS streamflow, soil moisture, etc). Because we compare spatial generalization, we used temporal validation splits for hyperparameter optimization. Complete details of hyperparameters, forcings, and static features for all four experiments are in Appendix C.
Model variants used.
Unless noted otherwise, pretrained encoders are frozen and only task heads are trained. We evaluate: (i) LSTM-SL: supervised LSTM baseline; (ii) StefaLand-direct: the StefaLand encoder together with its original decoder trained directly on the task data, without pretraining; (iii) StefaLand-resConn: pretrained encoder with residual pathway + LSTM decoder, our proposed architecture; (iv) StefaLand-noResConn: pretrained encoder with a simple adapter (no residual/LSTM); (v) StefaLand-scratch: same resConn head but StefaLand initialized randomly (no pretraining) and unfrozen (ablation test); (vi) TerraMind-resConn and PrithviWxC-resConn: EO/atmospheric foundation encoders frozen with the same residual head.
@@ 3.1 CAMELS Streamflow Prediction
To compare spatial generalization on a well benchmarked dataset, we follow (Feng et al., 2021), testing prediction in ungauged basins (PUB) and ungauged regions (PUR). These correspond to randomized spatial K-fold and regional-specific K-fold regimes, respectively. We use CAMELS (Addor et al., 2017, Newman et al., 2014), restricted to the 531-basin subset with clear watershed boundaries (Newman et al., 2017). Basins were divided into 10 random groups for PUB and 7 contiguous regions for PUR, employing leave-one-out in both cases. To avoid leakage, all CAMELS-overlapping stations were removed during pretraining for PUB, and entire regions were excluded for PUR.
Table 1: CAMELS Streamflow PUB and PUR Results
LSTM - SL 1.402 1.360 0.762 0.636 1.609 1.457 0.743 0.554
StefaLand - direct 1.882 1.849 0.538 0.395 1.982 1.949 0.230 0.201
StefaLand - resConn 1.111 1.068 0.869 0.717 1.344 1.334 0.801 0.635
StefaLand - no resConn 1.171 1.154 0.823 0.706 1.376 1.356 0.798 0.610
StefaLand - scratch 1.355 1.332 0.801 0.661 1.516 1.378 0.771 0.560
TerraMind - resConn 1.332 1.301 0.777 0.637 1.420 1.398 0.763 0.551
The foundation model pretraining clearly provides a rare boost to generalization capability compared to supervised learning models (LSTM-SL and StefaLand-direct) across both PUB and PUR (Table 1). StefaLand-resConn’s RMSE is almost $20\%$ lower than that of supervised LSTM in PUB and $17\%$ lower in PUR, while the correlation are noticeably higher. The value of pretraining is confirmed by StefaLand-scratch, which appears to have only minor advantage compared to LSTM. The residual-connection architecture enables effective integration of transformer features with temporal dynamics, showing a modest benefit compared to StefaLand-no resConn, but both are clearly stronger than either LSTM or StefaLand-direct. Finally, TerraMind-resConn has quite comparable metrics to LSTM; the pretrained TerraMind offers no extra generalization value in this case.
@@
We ran additional experiments that hybridize StefaLand with the HBV1.1 physics backbone on the same PUB/PUR splits, testing its ability to parameterize physics-based models. These hybrids achieved up to a 13% RMSE reduction and a 10% correlation gain compared to the LSTM–HBV1.1 baseline, demonstrating that pretraining strengthens hybridization with process models. By constraining predictions with physics while leveraging StefaLand features, these hybrids further improve upon the general results above and highlight the versatility of the approach. Full results are provided in Appendix B, Table B1.
@@
On a related note, the supervised LSTM is not an easy benchmark to surpass. The original multi-basin LSTM and subsequent large-scale comparisons (Kratzert et al., 2019, Feng et al., 2021) showed that vanilla Transformers generally fail to outperform LSTMs on rainfall–runoff prediction (Liu et al., 2024, Table 1 therein). The LSTM NSE values reported here are very similar to those in the domain literature (Feng et al., 2021). Broader evaluations likewise continue to rank LSTM-family models among the best-performing approaches (Lees et al., 2021), and even Google’s global flood-forecasting system adopts an encoder–decoder LSTM backbone (Nearing et al., 2024).
3.2 Global Streamflow
We designed a global-scale runoff prediction experiment to assess robustness and generalization worldwide. We filtered global basin datasets based on data completeness and retained 3,434 basins. To manage computational cost, we employed random hold-out sampling with three-fold cross-validation. Additionally, we implemented a regionally hold-out continental scenario (RH-C), excluding all basins from North America, South America, and Europe, respectively, from training, then evaluating on the excluded continent.
Table 2: Global streamflow prediction across 3,434 basins worldwide
Random holdout Regional holdout
(ungauged basins) (ungauged continents)
Model RMSE $\downarrow$ µbRMSE $\downarrow$ Corr $\uparrow$ RMSE $\downarrow$ µbRMSE $\downarrow$ Corr $\uparrow$
LSTM - SL 0.870 0.864 0.798 1.253 1.202 0.672
StefaLand - direct 0.749 0.751 0.843 1.075 1.048 0.697
TerraMind - resConn 1.156 1.111 0.580 1.234 1.158 0.670
Cross-continental transfer testing using three-fold validation. Detailed metric calculations in Appendix E.
Results indicate that StefaLand-direct outperformed both LSTM-SL and the fine-tuned vision-based Transformer model (TerraMind) across all metrics for either random or regional holdout (Table 2). Specifically, in the random holdout scenario, StefaLand achieved an RMSE of 0.749, approximately 14% lower than LSTM (0.87), with Corr improving to 0.843. Notably, StefaLand’s unbiased RMSE closely matched its RMSE, indicating errors were primarily random fluctuations rather than systematic bias. In contrast, TerraMind exhibited a higher RMSE of 1.156 and a low Corr of 0.580. We hypothesize this is because the pretraining satellite image data for the vision foundation network TerraMind did not have high relevance to hydrologic predictions like streamflow. We note that using TerraMind here was exploratory, as it was not specifically designed or pretrained for this use case.
Under the more challenging RH-C scenario, all models exhibited larger errors, but StefaLand continued to perform best. Its RMSE was 1.075, approximately 14.1% lower than LSTM’s 1.253, while also maintaining the highest Corr (0.697). Moreover, StefaLand’s ubRMSE remained close to its RMSE, confirming robust correction of regional-scale biases even under extreme out-of-domain conditions. Conversely, LSTM and TerraMind showed larger gaps between ubRMSE and RMSE, highlighting challenges in producing hydrologically-relevant features.
@@ 3.3 Global Soil Moisture
We evaluated finetuning StefaLand for soil moisture predictions following Liu et al. (2023a), using ISMN (Dorigo et al., 2011; 2013a). Even though there is a globally covering satellite-based product for soil moisture, the data quality can hardly match that of in-situ moisture sensors; thus the ability to generalize in-situ data is valuable. ISMN consists of 1,316 ground-based stations. We performed five-fold spatial cross-validation for random holdout and a regional holdout on Europe, training on all other continents while excluding European sites (129) for testing. We tested StefaLand-direct, StefaLand with and without residual connections (resConn / noResConn), a from-scratch ablation, an LSTM baseline, and a version using IBM’s TerraMind encoder with resConn. LSTM again serves as the established state-of-the-art baseline (Wang et al., 2024, Liu et al., 2023b).
@@
Table 3: Soil moisture prediction across 1,316 ISMN stations
Random location holdout Regional holdout
(random sites) (Europe)
Model RMSE $\downarrow$ µbRMSE $\downarrow$ Corr $\uparrow$ RMSE $\downarrow$ µbRMSE $\downarrow$ Corr $\uparrow$
LSTM - SL 0.073 0.055 0.764 0.112 0.053 0.510
StefaLand - direct 0.140 0.103 0.637 0.135 0.112 0.503
StefaLand - resConn 0.068 0.054 0.783 0.090 0.059 0.638
StefaLand - no resConn 0.075 0.057 0.741 0.095 0.058 0.545
StefaLand - scratch 0.074 0.058 0.749 0.108 0.064 0.528
TerraMind - resConn 0.083 0.062 0.694 0.101 0.080 0.519
PrithviWxC - resConn 0.081 0.060 0.703 0.103 0.079 0.523
5-fold spatial validation and cross-continental validation on Europe (129 sites). Detailed metric calculations in Appendix E.
Against these strong baselines, the soil moisture experiments confirm the superiority of the StefaLand-resConn architecture (Table 3). It achieves the best performance, with RMSE of 0.068 and correlation of 0.783 for random holdout, and maintains strong performance even in cross-continental testing on Europe (RMSE = 0.090, Corr = 0.638). The direct StefaLand model performs poorly because it lacks an effective mechanism to capture task-specific temporal dependencies. Similar to the streamflow cases above, TerraMind and PrithviWxC showed no performance benefit compared to LSTM, underscoring the challenge of repurposing vision foundation models across domains and the importance of pretraining with geoscience-relevant variables. The regional holdout on Europe further demonstrates StefaLand-resConn’s superior spatial generalization, achieving a 25% correlation improvement over the LSTM baseline in this difficult extrapolation scenario.
We want to emphasize that TerraMind and PrithviWxC were not designed for the hydrologic and land–surface prediction tasks studied here. They excel in their intended domains of Earth observation imagery and atmospheric variables, but are out-of-domain for dynamic land–surface and hydrologic modeling. We include them only to explore whether such models transfer any useful signal in our setting. These results should not be interpreted as evidence against their capability in other applications, but rather as a reflection of the mismatch between their pretraining objectives and the land–surface tasks we target. Nevertheless, we believe these benchmarks are helpful for clarifying their respective strengths, since the general AI community may not be familiar with these datasets and models that have been marketed as ‘any-to-any’ generative EO foundation models. Ultimately we view StefaLand as complementary to satellite-image foundation models, targeting a different but relevant data regime. We also could not identify other foundation models for land surface.
@@ 3.4 Soil Property Prediction
There are different soil datasets, each collected with different protocols and data processing techniques, resulting in significant discrepancies. In this test, we finetuned StefaLand to predict in-situ soil profile data from another dataset (ISRIC). This application can produce a seamless dataset that is consistent with a set of in-situ data, improving data availability and addressing systematic biases. In addition, it helps us understand the noise associated with each dataset. StefaLand’s pretraining soils dataset is HWSD, which has some overlap but also extensive differences from ISRIC, which is larger and potentially noisier. We finetuned StefaLand to predict one soil texture property (e.g., clay percentage) in ISRIC while masking the corresponding complementary attribute (e.g., sand) from the same profile to avoid information leakage, probing how easy it is to infer soil properties using other attributes such as climate, terrain and land cover. We compared StefaLand with and without pretraining against a supervised random forest baseline. Train-test splits can be found here C.2
Table 4: In-situ soil property prediction using ISRIC WoSIS data.
Model WoSIS Property Corr $\uparrow$ $R^{2}$ $\uparrow$
StefaLand - direct Clay 0.197 0.038
StefaLand - finetune Clay 0.509 0.259
Random Forest Clay 0.456 0.207
Linear Regression Clay 0.138 0.019
StefaLand - direct Sand 0.253 0.064
StefaLand - finetune Sand 0.704 0.495
Random Forest Sand 0.585 0.342
Linear Regression Sand 0.347 0.120
StefaLand with finetuning achieved markedly higher predictive power for both clay and sand fractions, substantially outperforming direct training, random forest, and linear regression baselines. The other models struggle in this task, with random forest (RF) scoring noticeably lower. This task is not dynamical prediction, but a test of the models ability to build deep representations of the landscape that can be adapted to infer certain attributes using cross-variable group connections. StefaLand-finetune’s advantage against random forest suggests there are indeed deep representations uncovered by pretraining but not exploited by random forest. Finetuning StefaLand is also apparently superior to training from scratch as supervised learning (StefaLand - direct), again highlighting the value of pretraining. These results highlight StefaLand’s ability to reconcile disparate attribute datasets and improve the utility of noisy in-situ observations.

### v3

@@ 3 Experiments
We tested the value of foundation model pretraining on 5 datasets and 6 experiments, including, streamflow on the CAMELS dataset on USA (a widely used benchmark dataset in hydrology), CAMELS streamflow prediction with hybrid model, global streamflow on Caravan, global in-situ soil moisture, global soil properties, and landslide susceptibility in Oregon, USA. For all experiments, hyperparameters were tuned with Ray Tune and kept consistent across model configurations within each experimental case (e.g., CAMELS streamflow, soil moisture, etc). Because we compare spatial generalization, we used temporal validation splits for hyperparameter optimization. Complete details of hyperparameters, forcings, and static features for all four experiments are in Appendix C.
Model variants and baselines.
We evaluate StefaLand alongside established baselines commonly used in hydrologic and geophysical time series modeling. As supervised baselines, we include LSTM-SL, which remains the dominant architecture for streamflow prediction and follows the same core model class used in prior large scale hydrology studies (Kratzert et al., 2019; Feng et al., 2021; Sabzipour et al., 2023). We additionally evaluate Informer (Zhou et al., 2021), Reformer (Kitaev et al., 2020), and DLinear (Zeng et al., 2022), which, although originally proposed for forecasting, are applied here in a sequence to sequence setting and are commonly included in recent hydrologic and geophysical modeling studies as representative transformer-based baselines.
To assess pretrained Earth representations, we include AlphaEarth LSTM and AlphaEarth ResConn, which feed pretrained AlphaEarth embeddings into an LSTM decoder with and without a residual task adapter, respectively (Brown et al., 2025). Because our task periods do not fully overlap AlphaEarth’s 2017–2024 temporal coverage, AlphaEarth embeddings are spatially aggregated over each basin and collapsed into a fixed set of 64 static features per site. For CAMELS streamflow and soil moisture, we additionally evaluate TabPFN, a tabular foundation model (Hollmann et al., 2025). Since TabPFN is non sequential and constrained by a fixed context window, we provide it with compact statistical summaries of each historical window while preserving identical data splits and evaluation protocols.
We also conducted an exploratory evaluation of additional Earth observation and atmospheric foundation models to assess their viability for land–surface prediction under matched adaptation strategies; these limited-scope results are reported in B.3
Our proposed method, StefaLand resConn (residual connection), combines a pretrained encoder with a residual adapter and LSTM decoder.
@@ 3.1 CAMELS Streamflow Prediction
To compare spatial generalization on a well benchmarked dataset, we follow (Feng et al., 2021), testing prediction in ungauged basins (PUB, random spatial K-fold) and ungauged regions (PUR, regional spatial K-fold). We use CAMELS (Addor et al., 2017; Newman et al., 2014), restricted to the 531-basin subset with clear watershed boundaries (Newman et al., 2017). Basins were divided into 10 random groups for PUB and 7 contiguous regions for PUR, employing leave-one-out in both cases. To avoid leakage, all CAMELS-overlapping stations were removed during pretraining for PUB, and entire regions were excluded for PUR. The success at these tasks would mean better flood forecasting information for populations in the world who do not have as many gauging stations around them.
Table 1: CAMELS streamflow performance under PUB evaluation (random spatial holdout; ungauged basins). Values are median $\pm$ standard error across folds.
Model RMSE $\downarrow$ ubRMSE $\downarrow$ Corr $\uparrow$ NSE $\uparrow$
LSTM SL 1.402 $\pm$ 0.04 1.360 $\pm$ 0.04 0.762 $\pm$ 0.01 0.636 $\pm$ 0.04
DLinear 2.012 $\pm$ 0.06 2.000 $\pm$ 0.06 0.598 $\pm$ 0.01 0.302 $\pm$ 0.48
Informer 2.262 $\pm$ 0.07 2.237 $\pm$ 0.08 0.521 $\pm$ 0.01 0.104 $\pm$ 0.08
Reformer 1.908 $\pm$ 0.09 1.871 $\pm$ 0.09 0.718 $\pm$ 0.00 0.270 $\pm$ 0.18
TabPFN 2.727 $\pm$ 0.09 2.725 $\pm$ 0.09 0.718 $\pm$ 0.01 0.509 $\pm$ 0.014
AlphaEarth LSTM 1.409 $\pm$ 0.10 1.393 $\pm$ 0.10 0.837 $\pm$ 0.11 0.618 $\pm$ 0.05
AlphaEarth resConn 1.361 $\pm$ 0.09 1.345 $\pm$ 0.09 0.859 $\pm$ 0.01 0.647 $\pm$ 0.04
StefaLand resConn 1.111 $\pm$ 0.04 1.068 $\pm$ 0.04 0.869 $\pm$ 0.01 0.717 $\pm$ 0.16
Table 2: CAMELS streamflow performance (PUR: regional spatial holdout; ungauged regions). Values are median $\pm$ standard error across folds.
Model RMSE $\downarrow$ ubRMSE $\downarrow$ Corr $\uparrow$ NSE $\uparrow$
LSTM SL 1.609 $\pm$ 0.24 1.457 $\pm$ 0.22 0.743 $\pm$ 0.02 0.554 $\pm$ 0.13
DLinear 2.019 $\pm$ 0.35 1.983 $\pm$ 0.34 0.597 $\pm$ 0.03 0.290 $\pm$ 0.99
Informer 2.332 $\pm$ 0.41 2.295 $\pm$ 0.37 0.497 $\pm$ 0.01 0.046 $\pm$ 0.27
Reformer 2.074 $\pm$ 0.33 1.978 $\pm$ 0.31 0.686 $\pm$ 0.02 0.257 $\pm$ 1.17
TabPFN 2.709 $\pm$ 1.08 2.501 $\pm$ 0.89 0.576 $\pm$ 0.11 0.328 $\pm$ 0.12
AlphaEarth LSTM 1.724 $\pm$ 0.85 1.685 $\pm$ 0.82 0.780 $\pm$ 0.07 0.456 $\pm$ 0.15
AlphaEarth resConn 1.727 $\pm$ 0.74 1.684 $\pm$ 0.71 0.790 $\pm$ 0.06 0.520 $\pm$ 0.10
StefaLand resConn 1.344 $\pm$ 0.21 1.334 $\pm$ 0.19 0.801 $\pm$ 0.02 0.635 $\pm$ 0.25
The results in Tables 1 and 2 demonstrate that foundation-model pretraining substantially improves spatial generalization relative to purely supervised approaches under both PUB and PUR evaluation. StefaLand-resConn achieves the strongest performance across all reported metrics, reducing RMSE by approximately 20% relative to the supervised LSTM baseline under PUB and by about 16 to 17% under PUR, while also yielding consistently higher correlation and NSE. Among the baselines, the supervised LSTM remains competitive and outperforms linear and transformer-based sequence models, while TabPFN performs poorly with substantially higher RMSE and lower NSE despite moderate correlation, indicating that a tabular formulation fails to capture key temporal structure in rainfall runoff dynamics. AlphaEarth-based variants provide notable improvements over the supervised LSTM, particularly in correlation, but remain well below StefaLand-resConn across all metrics, even with residual feature reuse.
@@
We ran additional experiments that hybridize StefaLand with the HBV1.1 physics backbone on the same PUB/PUR splits, testing its ability to parameterize physics-based models. These hybrids achieved up to a 13% RMSE reduction and a 10% correlation gain compared to the LSTM–HBV1.1 baseline, showing StefaLand’s . By constraining predictions with physics while leveraging StefaLand features, these hybrids further improve upon the general results above and highlight the versatility of the approach. Full results are provided in Appendix B, Table 9.
@@
On a related note, the supervised LSTM is not an easy benchmark to surpass. The original multi-basin LSTM and subsequent large-scale comparisons (Kratzert et al., 2019; Feng et al., 2021) showed that vanilla Transformers generally failed to outperform LSTMs on rainfall–runoff prediction and modified transforms essentially tied LSTM (Liu et al., 2024, Table 1 therein). The LSTM NSE values reported here are very similar to those in the domain literature (Feng et al., 2021). Google’s global flood-forecasting system adopts an encoder–decoder LSTM backbone (Nearing et al., 2024). The gains demonstrated are also substantial: to provide some context, when LSTM raised NSE from 0.64 to 0.73 (without ensemble), it was considered a generational change in predictive performance (Nearing et al., 2021; Feng et al., 2021), and no method other than ensembling more models cleanly surpassed default LSTM on the PUB test.
3.2 Caravan Global Streamflow
We designed a global-scale runoff prediction experiment to assess spatial robustness and generalization across diverse hydroclimatic regimes. We use the open-source Caravan dataset (Kratzert et al., 2023), a global community dataset that integrates meteorological forcings, catchment attributes, and observed streamflow for river basins worldwide. From the full Caravan archive (16,300 basins), we restricted our evaluation to the official CAMELS-family datasets, yielding 3,278 basins with relatively fuller streamflow records. From this set, we constructed a curated subset of 3,026 basins by excluding those with more than 75% missing streamflow observations during the 10-year evaluation period from 2010 to 2020, ensuring sufficient data quality for model training and testing. This filtering removes sparsely observed basins while preserving substantial geographic, climatic, and physiographic diversity. We employed a five-fold random spatial holdout protocol to evaluate generalization to unseen basins at the global scale.
Table 3: Caravan streamflow performance under random spatial holdout. Values are median $\pm$ standard error across folds
Model RMSE $\downarrow$ ubRMSE $\downarrow$ Corr $\uparrow$ NSE $\uparrow$
LSTM SL 1.623 $\pm$ 0.87 1.589 $\pm$ 0.84 0.496 $\pm$ 0.08 0.398 $\pm$ 0.74
DLinear 1.605 $\pm$ 0.99 1.453 $\pm$ 0.99 0.512 $\pm$ 0.07 0.412 $\pm$ 0.81
Informer 1.562 $\pm$ 1.04 1.471 $\pm$ 0.98 0.435 $\pm$ 0.07 0.256 $\pm$ 0.92
Reformer 1.574 $\pm$ 1.06 1.507 $\pm$ 0.97 0.451 $\pm$ 0.08 0.378 $\pm$ 0.87
AlphaEarth LSTM 1.533 $\pm$ 0.97 1.495 $\pm$ 0.89 0.523 $\pm$ 0.09 0.489 $\pm$ 0.89
AlphaEarth resConn 1.424 $\pm$ 1.03 1.406 $\pm$ 1.00 0.541 $\pm$ 0.07 0.514 $\pm$ 0.79
StefaLand resConn 1.457 $\pm$ 0.87 1.401 $\pm$ 0.83 0.589 $\pm$ 0.06 0.533 $\pm$ 0.70
Across the global Caravan benchmark (Table 3), StefaLand-resConn achieves the highest correlation and NSE among all evaluated models, indicating improved agreement with observed runoff dynamics under substantial spatial heterogeneity. While error-based metrics such as RMSE and ubRMSE vary considerably across basins, StefaLand-resConn consistently improves correlation relative to both supervised sequence models and alternative pretrained representations. AlphaEarth-based variants yield noticeable gains over purely supervised baselines, but remain below StefaLand-resConn across most metrics. The Caravan is a more noisy dataset due to global inconsistencies in data collection. These results suggest that attribute-centric pretraining combined with residual temporal adaptation provides a foundation to improve global hydrologic services.
@@ 3.3 Global Soil Moisture
We evaluated finetuning StefaLand for soil moisture predictions following (Liu et al., 2023a), using ISMN (Dorigo et al., 2011, 2013a). Even though there is a globally covering satellite-based product for soil moisture, the data quality can hardly match that of in-situ moisture sensors; thus the ability to generalize in-situ data is valuable. ISMN consists of 1,316 ground-based stations. We performed five-fold spatial cross-validation for random holdout and a regional holdout on Europe, training on all other continents while excluding European sites (129) for testing. LSTM again serves as the established state-of-the-art baseline (Wang and others, 2024; Liu et al., 2023b).
@@
Table 4: Soil moisture prediction across ISMN (random location holdout). Values are median $\pm$ standard error across folds.
Model RMSE $\downarrow$ ubRMSE $\downarrow$ Corr $\uparrow$
LSTM SL 0.073 $\pm$ 0.002 0.055 $\pm$ 0.001 0.764 $\pm$ 0.006
DLinear 0.088 $\pm$ 0.001 0.065 $\pm$ 0.001 0.612 $\pm$ 0.007
Informer 0.102 $\pm$ 0.002 0.082 $\pm$ 0.001 0.232 $\pm$ 0.012
Reformer 0.090 $\pm$ 0.002 0.071 $\pm$ 0.001 0.568 $\pm$ 0.015
TabPFN 0.068 $\pm$ 0.004 0.057 $\pm$ 0.003 0.404 $\pm$ 0.016
AlphaEarth LSTM 0.075 $\pm$ 0.001 0.062 $\pm$ 0.001 0.427 $\pm$ 0.019
AlphaEarth resConn 0.082 $\pm$ 0.001 0.067 $\pm$ 0.001 0.406 $\pm$ 0.012
StefaLand resConn 0.068 $\pm$ 0.001 0.054 $\pm$ 0.001 0.783 $\pm$ 0.005
Table 5: Soil moisture prediction across ISMN cross-continental validation on Europe (129 sites).
Model RMSE $\downarrow$ ubRMSE $\downarrow$ Corr $\uparrow$
LSTM SL 0.112 0.053 0.510
DLinear 0.093 0.051 0.623
Informer 0.088 0.063 0.358
Reformer 0.087 0.058 0.553
TabPFN 0.081 0.054 0.401
AlphaEarth LSTM 0.087 0.068 0.313
AlphaEarth resConn 0.090 0.071 0.308
StefaLand resConn 0.090 0.059 0.638
Against these baselines, the soil moisture experiments demonstrate advantages of the StefaLand-resConn architecture (Tables 4 and 5). For random spatial holdout, StefaLand-resConn attains the highest correlation (0.783) while also achieving the lowest ubRMSE (0.054) and matching the best RMSE (0.068), outperforming the LSTM baseline across all metrics. Although TabPFN achieves comparable RMSE, its substantially lower correlation indicates weaker agreement with temporal variability relative to sequence-based models.
The regional holdout on Europe represents a more challenging cross-continental generalization setting. In this case, StefaLand-resConn again achieves the highest correlation (0.638), exceeding all baselines including LSTM and linear or transformer-based models, while maintaining competitive RMSE and ubRMSE. In contrast, AlphaEarth-based variants show limited transfer performance, with both LSTM and residual-connection formulations yielding lower correlations despite similar error magnitudes.
Overall, these results highlight that residual adaptation on top of geoscience-pretrained representations is critical for robust soil moisture prediction, particularly under strong spatial distribution shift.
@@ 3.4 Soil Property Prediction
There are different soil datasets, each collected with different protocols and data processing techniques, resulting in significant discrepancies. In this test, we finetuned StefaLand to predict in situ soil profile data from another dataset (ISRIC). This application can produce a seamless dataset that is consistent with a set of in situ data, improving data availability and addressing systematic biases. In addition, it helps us understand the noise associated with each dataset. StefaLand’s pretraining soils dataset is HWSD, which has some overlap but also extensive differences from ISRIC, which is larger and potentially noisier. We finetuned StefaLand to predict one soil texture property (e.g., clay percentage) in ISRIC while masking the corresponding complementary attribute (e.g., sand) from the same profile to avoid information leakage, probing how easy it is to infer soil properties using other attributes such as climate, terrain, and land cover. We compared StefaLand finetuning against AlphaEarth features incorporated into a random forest model, as well as supervised random forest and linear regression baselines. Train test splits can be found here C.5.
Table 6: Soil composition prediction (clay). Values are median $\pm$ standard error across folds.
Model Corr $\uparrow$ $R^{2}$ $\uparrow$
Random Forest 0.456 $\pm$ 0.02 0.207 $\pm$ 0.01
Linear Regression 0.138 $\pm$ 0.01 0.019 $\pm$ 0.01
AlphaEarth RF 0.462 $\pm$ 0.03 0.205 $\pm$ 0.02
StefaLand finetuned 0.509 $\pm$ 0.01 0.259 $\pm$ 0.01
Table 7: Soil composition prediction (sand). Values are median $\pm$ standard error across folds.
Model Corr $\uparrow$ $R^{2}$ $\uparrow$
Random Forest 0.585 $\pm$ 0.01 0.342 $\pm$ 0.01
Linear Regression 0.347 $\pm$ 0.01 0.120 $\pm$ 0.01
AlphaEarth RF 0.615 $\pm$ 0.01 0.373 $\pm$ 0.01
StefaLand finetuned 0.704 $\pm$ 0.01 0.495 $\pm$ 0.01
StefaLand finetuning achieves the strongest performance for both clay and sand prediction, outperforming linear regression, random forest, and AlphaEarth RF baselines in terms of both correlation and $R^{2}$. While AlphaEarth RF improves over standard random forest baselines, it remains consistently below StefaLand finetuned across both soil properties. These results indicate that StefaLand’s pretrained representations are particularly effective when adapted to infer soil texture attributes from complementary variables and static environmental context.